40 A positive point charge is located at the origin of an x-y coordinate system, and an electron is placed at a location where the electric field due to the point charge is given by E Eo(i +j),where Eo is positive. What is the direction of the force on the electron?
Answer #1

Here ,

electric field is E = Eo * (i + j)

for the electric force on electron

Force = q* E

Force = -e * Eo * (i + j)

direction of force = - (i + j)

direction of force = 180 + tan(1/1)

direction of force = 225 degree

the direction of force is 225 degree with positive x axis